### v1

# Towards AI-Driven Policing: Interdisciplinary Knowledge Discovery from Police Body-Worn Camera Footage

Abstract
This paper proposes a novel interdisciplinary framework for analyzing police body-worn camera (BWC) footage from the Rochester Police Department (RPD) using advanced artificial intelligence (AI) and statistical machine learning (ML) techniques. Our goal is to detect, classify, and analyze patterns of interaction between police officers and civilians to identify key behavioral dynamics, such as respect, disrespect, escalation, and de-escalation. We apply multimodal data analysis by integrating video, audio, and natural language processing (NLP) techniques to extract meaningful insights from BWC footage. We present our methodology, computational techniques, and findings, outlining a practical approach for law enforcement while advancing the frontiers of knowledge discovery from police BWC data.
Index Terms:
Body-worn cameras, Multimodal Data Analysis, Audio Processing, Speaker Diarization, Natural Language Processing, Police Training
I Introduction
Police body-worn cameras (BWCs) are expanding video-gathering tools used for evidence collection that allow researchers, crime analysts, and academics to extract insights into police dynamics in real-time. The extensive volume of data generated by BWC footage presents significant challenges across the criminal justice system, complicating efforts to analyze this information effectively and assess police-civilian interactions [1]. To address these challenges, this paper introduces the OpenBWC (Open Body-Worn Camera) framework, an open-source, multimodal system designed to leverage artificial intelligence (AI) for analyzing patterns of behavior captured in BWC footage.
     Using open-source AI technologies in a responsible and unbiased manner, this framework aims to support informed decision-making in policing practices. Our approach integrates advanced speaker separation, audio transcription, and natural language processing (NLP) techniques to examine police-civilian interactions, enabling systematic analysis and classification of behavioral patterns. By employing these methods, we aim to first classify police-civilian encounters based on key elements such as de-escalation, professionalism, and conflict resolution. Establishing this classification framework will allow us to identify critical factors influencing these outcomes. The goal is to translate analytical findings into practical strategies to improve police performance, with potential training approaches informed by collaboration with the Rochester Police Department (RPD). Current efforts focus on developing reliable methods to classify police-civilian encounters based on key behavioral and contextual elements. Consequently, this study investigates how multimodal AI techniques can systematically analyze and classify behavioral patterns in BWC footage, providing a structured foundation to improve police training, police-civilian interactions, and everyday policing practices. This approach also considers the impact of civilian behavior or error, aiming to advance public safety, equity, and the broader pursuit of justice.
II Related Work
Researchers have used police BWC footage as data to learn about policing for several years already (Makin et al., 2018) [2]. With advancements in AI and NLP, this research has evolved to support large-scale analysis of recorded interactions. Computational methods have examined officer behavior, communication strategies, and officer-citizen dynamics, offering valuable contributions to policing research.
     Camp and Voigt (2025) [3] proposed a framework that applies NLP techniques to BWC footage to identify behavioral patterns, assess communication styles, and evaluate policing practices. Their approach emphasizes scaling up analysis to highlight conversational themes, officer respect, and situational factors contributing to escalation or de-escalation. While they describe how supervised and unsupervised NLP models could extract linguistic features such as politeness, apology, and reassurance, their work remains largely theoretical, with the framework’s practical implementation and performance yet to be demonstrated. While their framework shows a potential for automating behavioral analysis, the authors note that its accuracy may depend on factors such as the quality of transcriptions, the reliability of the model, and the complexity of conversations.
Figure 1: OpenBWC System Workflow: Multimodal Data Processing and Analysis Pipeline.
Additionally, they point out that noisy environments and different ways of speaking may create challenges, highlighting the importance of thorough testing in real-world policing situations. This framework’s ability to automate analysis has the potential to illuminate the nature of interactions that are often challenging to observe manually.
     As a related initiative, Camp et al. (2024) [4] used BWC footage to assess the effects of communication-based training programs designed to improve officer behavior during traffic stops. Their study employed NLP models to analyze features such as respectful language, tone shifts, and conversational structure. Results indicated improved officer behavior and increased professionalism following targeted interventions, emphasizing the value of AI language models for training evaluation.
     Researchers have developed innovative annotation and labeling strategies to improve the scalability and efficiency of BWC footage analysis. One such tool is the Computer Vision Annotation Tool-Body-Worn Video (CVAT-BWV), a web-based platform designed specifically for annotating BWC footage [5]. CVAT-BWV simplifies labeling with user-friendly interfaces for identifying key behaviors, conversational cues, and non-verbal gestures. By combining automatic labeling techniques with manual annotations, this system enhances the accuracy of analyzing police-civilian interactions and identifying patterns in officer conduct.
     Moreover, recent efforts to improve the analysis of BWC footage have led to the development of video-based systematic social observation (VBSSO) methods. McCluskey et al. (2023) introduced VBSSO as a structured approach for studying police-civilian interactions using video data from police archives [6]. Their method includes strategies for sampling relevant BWC files by linking them to a department’s record management system (RMS), allowing for more efficient data selection. These techniques have influenced our framework, improving its ability to analyze behavioral patterns in police-civilian interactions.
     Building on these foundations, we combine open-source speaker separation models like SepReformer [7] and transcription systems like WhisperAI [8] to assess the reliability of various tools in handling the challenging, noise-filled environments commonly found in BWC footage. Additionally, transcriptions were summarized using the open-source Large Language Model (LLM) Llama 3.3 model [9]. By integrating these tools with the open-source database PostgreSQL for structured data organization, our system ensures scalable and reliable data retrieval to support comprehensive pattern analysis. This open-source design promotes accessibility and reproducibility, enabling researchers, machine learning (ML) practitioners, and police departments to adopt and adapt the framework for their own investigations.
III Methodology
Our multimodal framework, as illustrated in Figure 1, is designed to analyze audio, text, and image data in an integrated manner. To combine the strengths of each modality, the system applies an ensemble learning approach, described in III-A, that integrates audio, text, and visual features into a unified analysis. This design uses both quantitative and qualitative techniques to analyze behavioral patterns, conversational dynamics, and environmental context. The chosen methods emphasize data segmentation, advanced ML models, and structured data organization to ensure scalability and accuracy.
     At the core of the OpenBWC system is the knowledge extraction algorithm (Algorithm 1), which provides a step-by-step procedure for processing complex BWC footage. By leveraging advanced audio processing, transcription models, and NLP techniques, this algorithm extracts meaningful observations at each stage. In parallel, knowledge extraction runs alongside every pipeline component to enrich the data with semantic labels and additional insights.
Data. The dataset used in this research comprises 834 FOIL (Freedom of Information Law) videos obtained from publicly available sources through FOIL requests. These recordings capture a broad spectrum of interaction contexts, including police-civilian conversations, emergency responses, and high-pressure situations. Each BWC recording, stored as MP4 files, was systematically extracted from storage systems for further processing, including audio enhancement, transcription, and structured data organization to enable systematic and thorough analysis.
TABLE I: Summary of BWC FOIL Video Dataset by Year.
Year No. of Videos Avg. Length (min)
2017 21 50.95
2019 43 5.78
2022 219 16.72
2023 365 13.62
2024 186 41.34
Total 834 128.41
Table I summarizes the distribution of videos by year, presenting the number of recordings and their average lengths. The data exhibit substantial variation, with 2024 recordings showing the longest average durations and mid-period years reflecting shorter interactions. This spread reflects a wide range of police-civilian interactions over time, offering a strong basis for analyzing contextual trends in BWC footage.
Audio Processing Pipeline. The audio pipeline was designed to handle long and complex recordings by applying chunking, speaker separation, and transcription models.
Algorithm 1 Knowledge Extraction from BWC Footage
1:procedure AnalyzeFootage(videoDataset)
2:    for each video in videoDataset do
3:         audio $\leftarrow$ ExtractAudio(video)
4:         videoChunks $\leftarrow$ SplitVideo(video, time)
5:         for each chunk in videoChunks do
6:             speakers $\leftarrow$ SourceSeparation(chunk.audio)
7:             for each speaker in speakers do
8:                  transcript[speaker] $\leftarrow$ TranscribeAudio( speaker.audio)
9:             end for
10:              mergedTranscript $\leftarrow$ MergeTranscripts(transcript)
11:         end for
12:          fullTranscript $\leftarrow$ SummarizeTranscripts(mergedTranscript)
13:         VerifyAndCorrect(fullTranscript)
14:         insights $\leftarrow$ NLPAnalysis(fullTranscript)
15:         SaveInsights(insights)
16:    end for
17:end procedure
The steps of this process are outlined in Algorithm 1, which complements the description provided in this section. Following the algorithm alongside the narrative may aid in understanding the sequence of operations more clearly. In Line 1, the algorithm begins by iterating through the dataset of BWC videos. Each video undergoes audio extraction (Line 2), followed by segmentation into 15-second chunks (Line 3) during initial experiments. While this approach improved system responsiveness and reduced GPU load, it introduced challenges with conversational continuity and keeping track of context. Long audio files made the system work harder and slowed down transcription. To address this, we switched to 30-second chunks, which helped conversations sound more natural. This choice reflected a trade-off: shorter chunks often broke conversations into smaller parts, making it harder to maintain context and track speakers. In contrast, longer chunks better preserved conversational flow, but required more computing power and made it harder to separate speakers during source separation. However, in cases with fewer speakers or access to more powerful hardware, longer chunks may still be a suitable option. By working with these manageable audio units, the system could better capture conversational turns, improving both speaker identification and transcription quality.
     Following segmentation (Lines 4-10), each audio chunk undergoes source separation using the SepReformer model—an asymmetric encoder-decoder architecture designed for noisy environments, as illustrated in Figure 2. SepReformer is based on Blind Source Separation (BSS) principles, a technique that separates individual speaker signals from a mixture of overlapping audio sources without prior knowledge of the speakers. BSS models are particularly effective in noisy environments where traditional separation methods struggle. Inspired by Bando et al. (2023) [10], our system uses SepReformer’s advanced encoding-decoding structure to isolate individual speakers with improved accuracy. In this process, each audio chunk (e.g., chunk_1.wav, chunk_2.wav) is first encoded through an Audio Encoder, which extracts key audio features in the form of compact representations of the input sound that capture important patterns such as tone, pitch, and energy. These features, also called latent sequence features, help distinguish and separate speakers in a mixture. The encoded features are then processed by the Source Separation Module, which isolates individual speakers from the mixed audio. Finally, the separated audio streams are decoded through the Audio Decoder to produce distinct waveforms corresponding to each speaker. This method generates audio outputs for each participant, improving transcription accuracy and enabling more precise dialogue analysis in multi-party conversations.
Figure 2: Multi-Speaker Audio Processing Flow.
Once speaker separation was complete, transcribed segments were merged into a full transcript for each video (Line 12). This step used WhisperAI (base model), optimized for real-world audio challenges like noise, overlapping speech, and varied accents—common in BWC footage.
     In this context, accuracy goes beyond word-for-word correctness. It includes whether the transcription makes the conversation understandable and preserves the speaker’s emotional tone, timing, and intent, particularly in high-pressure moments. We observed that performance tended to be more reliable in routine interactions, such as traffic stops, where speech patterns are more predictable and audio is generally clearer.
     Based on our observations, challenging videos—such as those with chaos, shouting, or loud background noise—were harder to transcribe accurately. Common issues included repeated phrases, incorrect speaker labels, or missed tone shifts. These moments are especially important for evaluating system performance in high-stakes interactions like de-escalation.
Text Analysis Pipeline. The text analysis stage builds upon the transcribed audio by identifying conversational themes and structuring the data for retrieval.
     To ensure accurate speaker attribution, each transcript was aligned with the speaker-separated audio, which allowed us to determine whether the speaker was a police officer or a civilian. As part of our review, we randomly examined a subset of videos focused on routine traffic stops and standard police checks, where officers were typically the dominant speakers—often asking procedural questions, giving instructions, and assisting civilians.
     In Line 13 of Algorithm 1, the full transcript is checked for alignment with the separated audio and metadata to verify speaker roles and maintain dialogue traceability. Once verified, Line 14 applies advanced NLP models—such as Llama 3.3—to extract meaningful observations from each interaction, including indicators of de-escalation, tone, and respect.
Database and Retrieval System. To enable scalable data storage and efficient search, the processed results were organized in a PostgreSQL database. This structured database was designed to support rapid data retrieval and facilitate comprehensive pattern analysis. One of the key features of the database is its indexed summaries, which significantly enhance search speed and filtering efficiency. By indexing key details such as speaker identifiers and conversational themes, the system ensures that relevant data can be accessed quickly and accurately.
     The database design was also built with scalability in mind. As the OpenBWC system continues to expand, this architecture can accommodate additional data sources, ensuring the system remains flexible and efficient as new records are integrated.
III-A Ensemble Learning for Multimodal Analysis
To achieve comprehensive knowledge extraction from the BWC data, the OpenBWC system employs a mathematically driven ensemble learning framework designed to integrate diverse data features. By combining audio, text, and visual information, this method improves prediction accuracy.
     Let $\hat{f}_{\text{Audio}}$, $\hat{f}_{\text{Text}}$, and $\hat{f}_{\text{Image}}$ represent feature functions derived from audio, text, and image data, respectively. Each of these functions is defined as follows:
$\hat{f}_{\text{Audio}}(\text{BWC Video})=h_{L}\circ h_{L-1}\circ\cdots\circ h_{2}\circ h_{1}(x)$
Where:
  • $x$ is the original audio waveform from BWC footage. The length of the audio can impact processing—longer recordings (e.g., over 30 minutes) often led to speaker separation errors in our experiments, where the system began to confuse or switch speakers later in the interaction, especially during complex or overlapping conversations.
  • Each transformation layer $h_{i}$ represents a feature extraction stage applied to the audio data, such as speaker diarization.
  • The final transformation layer $h_{L}$ extracts high-level acoustic features, such as speaker-specific voice patterns.
For the text analysis pipeline, the system extracts linguistic features from transcriptions using the following transformation function:
$\hat{f}_{\text{Text}}(\text{BWC Video})=\text{Llama 3.3}\circ\text{WhisperAI}(x)$
Where:
  • $x$ is the transcribed text produced by WhisperAI.
  • WhisperAI handles noise-robust transcription, while Llama 3.3 performs deep semantic analysis by identifying language markers such as politeness, de-escalation attempts, and reassurance cues.
For visual analysis, the system defines:
$\hat{f}_{\text{Image}}(\text{BWC Video})=v_{L}\circ v_{L-1}\circ\cdots\circ v_{2}\circ v_{1}(y)$
Where:
  • $y$ is the sequence of video frames extracted from the BWC footage.
  • Each layer $v_{l}$ represents a visual processing transformation, such as object detection and motion tracking.
The final ensemble model combines these three feature representations to create a unified model. The integrated function is defined as:
$\hat{g}(\text{BWC Video})=\alpha\hat{f}_{\text{Audio}}+\beta\hat{f}_{\text{Text}}+\gamma\hat{f}_{\text{Image}}$
Where:
  • $\alpha$, $\beta$, and $\gamma$ are weighting coefficients that control the contribution of each data modality.
  • The weighting values are calibrated during training to maximize the model’s accuracy.
This ensemble model allows the OpenBWC system to exploit the complementary strengths of each data type—acoustic cues for speaker identification, textual information for dialogue analysis, and visual data for environmental context.
IV Evaluation
The selected tools—particularly WhisperAI and SepReformer—were tested on challenging BWC footage containing background noise, overlapping speech, and varied speaker dynamics. To assess transcription quality, we manually reviewed a subset of BWC footage from routine traffic stops and compared the output transcripts with the separated audio streams. This review focused on evaluating both the accuracy of individual phrases and the clarity of speaker attribution. We found that the majority of the transcribed content—particularly officer speech—was reliably captured. This aligns with expectations, as the officer’s proximity to the BWC microphone often results in clearer audio input.
     Nonetheless, we noted instances of minor issues, including repeated phrases, occasional misinterpretations, and contextually inconsistent segments—particularly during moments of overlapping speech or increased background noise. Despite these limitations, the tools demonstrated strong baseline performance in structured interactions such as routine traffic stops.
     To further support interpretability, we applied the Llama 3.3 language model to summarize cleaned transcripts of police-civilian interactions. We guided the model to append relevant tags at the end of each summary, specifically indicating the presence of use of force, escalation, or violence. This helped us quickly identify key themes while keeping the context of each event. The model also merged the separate transcripts into one clear, speaker-labeled version, which made it easier to follow what happened. While these outputs are not a replacement for human judgment, they offer a practical starting point for downstream analysis, review, and tagging.
     The integration of WhisperAI, SepReformer, and Llama 3.3 with a structured data storage system not only supports current analysis but also lays the groundwork for future evaluations, refinements, and replicable studies within the OpenBWC framework.
V Limitations
Several challenges affected the system’s performance and accuracy during development.
  1. 1.
    GPU Memory Constraints. Processing long audio recordings, especially those that span several hours, requires substantial GPU memory. The SepReformer model required extensive resources for such segments. To address this, the recordings were divided into 30-second chunks. RIT’s Research Computing Services [11] provided high-performance GPUs, enhancing processing speed and enabling efficient handling of large datasets.
  2. 2.
    Noise Distortion. BWC footage often contains significant background noise, including sirens, vehicle sounds, and overlapping speech. Although noise reduction methods made audio clearer, sometimes too much filtering removed important details like how a person sounded or subtle cues in their voice. While text transcriptions capture the words spoken, vocal details like tone and intonation carry important meaning, emotions, and intentions—details that are often lost with too much audio filtering. Finding the right balance between reducing noise and keeping these vocal details is essential for accurately understanding noisy police encounters. Additionally, the officer’s voice is often captured with greater clarity than that of civilians, likely due to the proximity of the BWC’s microphone and its orientation toward the officer’s speech. This asymmetry in audio capture can reduce the transcription quality for civilian speech in dynamic or noisy settings.
  3. 3.
    Speaker Overlap. Overlapping speech in BWC recordings complicated transcription accuracy. Although the SepReformer model effectively separated speaker audio when only two speakers were present, it struggled in chaotic scenarios characterized by rapid interruptions or when three or more speakers overlapped. This issue impacts transcription quality during precisely the types of incidents that police managers prioritize for review—high-stakes, complex interactions. This finding aligns with previous research, such as Willits and Makin (2017) [12], who reported that most of the use-of-force cases recorded by BWCs lacked sufficient audio and video clarity for human coders to reliably interpret events. Improving audio separation techniques for multi-speaker situations represents an important area for future work, given their significance in understanding police-citizen encounters.
These technical limitations are particularly concerning in the context of public oversight and accountability, when clear and reliable audio is most important, like during high-stress or intense moments. Future work will aim to evaluate where, when, and why accuracy breaks down to inform both technical improvements and policy-level decisions about the appropriate use of BWC-derived evidence.
VI Future Work
Future work will expand the multi-modal approach to include voice stress detection, sentiment analysis, and object detection to identify emotionally charged segments for more targeted manual review in high-stakes scenarios. Improved diarization models may also enhance speaker identification accuracy in complex interactions. We also plan to develop cross-referencing capabilities that allow for linking multiple incidents, enhancing pattern detection, and enabling comparison across recordings of the same event captured from different BWCs. This will help address challenges posed by concatenated video releases and support a more accurate, multi-perspective understanding of dynamic police-civilian encounters. Collaboration with police officers will further support the interpretation of algorithm-clustered incidents by grounding findings in field expertise.
VII Conclusion
We introduced a framework that demonstrates the potential of combining AI-driven audio, text, and image processing with advanced AI techniques for knowledge discovery from BWC footage. Analyzing repositories of BWC footage does not require access to proprietary software; instead, the tools can be open source, and the resulting algorithms, built to analyze publicly owned data, can also remain public. This promotes transparency, reproducibility, and community engagement in future research. Building on the current analysis of a subset of approximately 834 videos, we identified key gaps—particularly in reliably assigning speaker identity during overlapping speech and maintaining transcription accuracy in high-noise, high-interruption scenarios. These limitations introduce uncertainty in the transcribed content. Addressing these gaps will be central to improving confidence in the outputs.
Acknowledgment
This research is supported by a grant from the U.S. Department of Justice (DOJ) through its Office of Justice Programs (OJP) and Bureau of Justice Assistance (BJA), awarded to the Rochester Police Department (RPD).